OpenStudy (anonymous):

Hi! Can anyone explain to me how to calculate log(x) without a calculator? For example log81 and log36 ?? I just learned logarithms and familiar with how it is written using bases but there is nothing that explains this to me. Also how does 3^x = 81 become xlog3 = log81 ??

OpenStudy (anonymous):

hmm ok i just found that the base is 10 for log(x)

OpenStudy (mathstudent55):

To answer your second question first, here's a rule of logs: \( \log x^n = n \log x\) Also, if you have an equation, you can take the log of both sides. Start with \(3^x = 81\) First, take the log of each side: \( \log 3^x = \log 81\) Use the rule above: \(x \log 3 = \log 81\) We can go further, if you realize that \( 81 = 3^4\) \( x \log 3 = \log 3^4\) \(x \log 3 = 4 \log 3\) \(x = \dfrac{4 \log 3}{\log 3} \) \(x = 4\)

OpenStudy (anonymous):

ok thanks for some reason that rule is not shown in my module, also does not state that log(x) uses a common base of 10.... -.- thanks!!